# floq: Training Critics via Flow-Matching for Scaling Compute in Value-Based RL

### v1

@@ 1 Introduction
Motivated by these empirical results, in this paper, we ask: can iterative computation also improve value estimation in reinforcement learning (RL)? Specifically, we are interested in improving the estimation of the Q-value function. While Q-functions map state-action inputs to a scalar value, they are known to be highly complex and difficult to fit accurately (e.g., [10]). Standard temporal-difference (TD) learning used to train Q-functions struggles to leverage capacity of deep networks [38, 39, 5, 50, 21], often resulting in poor generalization. These problems are further exacerbated in the offline RL problem setting [45, 36], where we must learn entirely from static datasets. This motivates exploring architectures that spend compute iteratively to estimate value functions, potentially yielding more accurate Q-values, and as a result, better downstream policies.
@@ 3 Background and Preliminaries
$\!\!\!L(\theta)=\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D},a^{\prime}\sim\pi(\cdot|s^{\prime})}\left[\left(r(s,a)+\gamma\bar{Q}(s^{\prime},a^{\prime})-Q_{\theta}(s,a)\right)^{2}\right],$ (3.1)
@@
$\frac{d}{dt}\psi_{\theta}(t,\mathbf{x})=v_{\theta}(t,\psi_{\theta}(t,\mathbf{x})),\quad\psi_{\theta}(0,\mathbf{x}(0))=\mathbf{x}(0)$ (3.2)
@@
$\min_{\theta}\;\mathbb{E}_{\mathbf{x}(0),\mathbf{x}(1),t}\left[\left\|v_{\theta}(t,\mathbf{x}(t))-\frac{(\mathbf{x}(1)-\mathbf{x}(0))}{1-0}\right\|_{2}^{2}\right].$ (3.3)
@@
In contrast to standard Q-functions parameterized by deep neural networks that map state-action pairs to scalar values, floq parameterizes a time-dependent, state-action-conditioned velocity field $v_{\theta}(t,\mathbf{z}\mid s,a)$ over a one-dimensional latent input $\mathbf{z}\in\mathbb{R}$. At $t=0$, this input $\mathbf{z}$ is sampled from the uniform distribution $\mathrm{Unif}\leavevmode\nobreak\ [l,u]$, where $l$ and $u$ are scalars that define the range of initial sample noise used for training. The velocity field transforms the initial sample $\mathbf{z}$ into a distribution over the Q-value. We will train floq such that the learned distribution of Q-values match a Dirac-Delta around the groundtruth Q-function, i.e., $\psi_{\theta}(1,\mathbf{z}|s,a)\sim\delta_{Q^{\pi}(s,a)}$ at $t=1$. We can obtain the Q-value sample by numerically integrating the ODE using the Euler method. One instantiation is shown below. $\forall j\leq K$:
$\psi_{\theta}(\nicefrac{{j}}{{K}},\mathbf{z}\mid s,a)=\mathbf{z}+\frac{1}{K}\sum_{i=1}^{j}v_{\theta}\left(\frac{i}{K},\psi_{\theta}\left(\nicefrac{{i-1}}{{K}},\mathbf{z}\mid s,a\right)\Big{|}s,a\right).\leavevmode\nobreak\ \leavevmode\nobreak\ Q(s,a,\mathbf{z}):=\psi_{\theta}(1,\mathbf{z}\mid s,a)$ (4.1)
@@ 4.2 Training Loss for the floq Architecture
$\mathcal{L}_{\texttt{{floq}}{}}(\theta)=\mathbb{E}_{\mathbf{z},t}\left[\left\|v_{\theta}(t,\mathbf{z}(t)\mid s,a)-\frac{({y}(s,a)-\mathbf{z})}{1-0}\right\|_{2}^{2}\right].$ (4.2)
@@ 4.3 Preventing Flow Collapse: How to Make floq Work Well?
Design choice 1: Distribution of the initial noise sample. As shown in prior works [47, 48], rescaling the source noise leaves the target distribution unchanged, but it alters the curvature of the transport trajectories. Interestingly, this effect seems to be particularly pronounced when applying flow-matching to scalar TD-learning (see Figure 12 in experiments). As such, we find that that setting the bounds $l$ and $u$ for the distribution of the initial noise, $\mathrm{Unif}\leavevmode\nobreak\ [l,u]$ greatly affects the performance of floq.
@@
Summary: floq architecture and training floq parameterizes the Q-function via a learned velocity field. We train this velocity field with a linear flow-matching loss against the target Q-value (Eq. 4.2). To ensure that floq scales capacity, we carefully choose the distribution of the initial noise $\mathrm{Unif}\leavevmode\nobreak\ [l,u]$ to be wide and overlapping with range of target Q-values. We encode the interpolant using a categorical representation. We represent the input $t$ to $v_{\theta}$ using a Fourier encoding. See Algorithm 1 for pseudocode and details.
@@ 5.1 Main Offline RL Results
Comparisons and evaluation protocol. In addition to a floq critic, we build our experiments to utilize flow-matching policies. This implies that flow-Q learning (FQL[59], which utilizes a flow-matching policy but with a monolithic Q-network, is our main comparison. The FQL work reports results with 11M training steps; for a fair comparison, we additionally re-run FQL for 22M steps and report both sets of results. We run floq with a default set of hyperparameters across tasks, but on the more challenging humanoidmaze-large and antmaze-giant tasks we found a larger batch size of $512$ to be more effective, which we use for both FQL and floq on these environments. Beyond FQL, we compare against three recent state-of-the-art offline RL algorithms, all of which rely on monolithic Q-functions: (i) ReBRAC [71], which corresponds to the strongest-performing method with a monolithic Q-network and a Gaussian policy, and thus provides a competitive non-flow baseline; (ii) DSRL [73] that adapts a diffusion-based behavior cloning policy by performing RL over its latent noise space, improving over FQL; ; and (iii) SORL [11] that leverages shortcut flow models to improve the training and inference scalability of flow policies. Note that none of them utilize a flow-matching Q-function, but do innovate across various properties of policy training. Comparing to the strongest methods that innovate on the design of the policy allows us to evaluate the importance of and benefits from utilizing flows for a critic.
@@
Since DSRL [73] only evaluates on the 10 default tasks (one default task per environment), we also provide an additional results table on only the default tasks in each environment in the appendix (Table 3). On the default tasks, we find that floq outperforms DSRL (20% for DSRL vs. 45% for floq), improving by over $2\times$ in success rate. These results establish the efficacy of floq, and generally, performance improvements in offline RL from scaling Q-functions via flow-matching models.
@@ Acknowledgements
We thank Max Sobol Mark, Zheyuan Hu, Aravind Venugopal, Seohong Park, Mitsuhiko Nakamoto, and Amrith Setlur for feedback on an earlier version of the paper and informative discussions. We thank members of the CMU AIRe lab for feedback and support. AK thanks Max Simchowitz for discussions on flow-matching and diffusion models. This work was supported by a Schmidt Sciences AI2050 fellowship and the Office of Naval Research under N00014-24-12206. We thank the TRC program of Google Cloud and NCSA Delta for providing computational resources that supported this work.
@@ Appendix A Additional Results for floq
antsoccer-arena-task2 22 $\pm 2$ 0 $\pm 1$ 96 $88$ $\pm 3$ $89$ $\pm 4$ $97$ $\pm 2$ $97$ $\pm 2$
@@
Table 3: Offline RL results (Default Tasks). floq achieves competitive or superior performance compared to the baselines. “Hard” tasks refers to the set of default tasks where the FQL baseline score below 50% performance. floq is especially more performant on these hard tasks over prior comparisons, where its performance more than doubles over baseline FQL.
@@
Env (Default Task) BC ReBRAC DSRL SORL FQL (1M) FQL(2M) floq (Def.) floq (Best)
antmaze-large 0 $\pm 0$ $91$ $\pm 10$ $40$ $\pm 29$ $93$ $\pm 2$ $80$ $\pm 8$ $85$ $\pm 4$ $94$ $\pm 4$ $94$ $\pm 4$
antmaze-giant 0 $\pm 0$ $27$ $\pm 22$ 0 $\pm 0$ $12$ $\pm 6$ $11$ $\pm 16$ $14$ $\pm 29$ $70$ $\pm 8$ $86$ $\pm 4$
hmmaze-medium 11 $\pm 0$ $16$ $\pm 9$ $34$ $\pm 20$ $67$ $\pm 4$ 1919 $\pm 12$ $58$ $\pm 25$ $98$ $\pm 1$ $98$ $\pm 1$
hmmaze-large 0 $\pm 0$ 22 $\pm 1$ $10$ $\pm 12$ 2020 $\pm 9$ 88 $\pm 5$ $14$ $\pm 10$ $52$ $\pm 8$ $52$ $\pm 8$
antsoccer-arena 11 $\pm 0$ 0 $\pm 0$ $28$ $\pm 0$ $54$ $\pm 5$ $39$ $\pm 6$ $49$ $\pm 11$ $49$ $\pm 10$ $49$ $\pm 10$
cube-single 33 $\pm 1$ $92$ $\pm 4$ $93$ $\pm 14$ $99$ $\pm 0$ $96$ $\pm 1$ $94$ $\pm 5$ $99$ $\pm 2$ $99$ $\pm 2$
cube-double 0 $\pm 0$ 77 $\pm 3$ $53$ $\pm 14$ $33$ $\pm 8$ $36$ $\pm 6$ $29$ $\pm 8$ $72$ $\pm 15$ $72$ $\pm 15$
scene 11 $\pm 1$ $50$ $\pm 13$ $88$ $\pm 9$ $89$ $\pm 9$ $76$ $\pm 9$ $78$ $\pm 7$ $83$ $\pm 10$ $83$ $\pm 10$
puzzle-3x3 11 $\pm 1$ 22 $\pm 1$ 0 $\pm 0$ - $\pm$- $16$ $\pm 5$ $14$ $\pm 4$ $17$ $\pm 6$ $17$ $\pm 6$
puzzle-4x4 0 $\pm 0$ $10$ $\pm 3$ $37$ $\pm 13$ - $\pm$- $11$ $\pm 3$ 55 $\pm 2$ $12$ $\pm 4$ 1919 $\pm 5$
Average Score (All Tasks) 11 $30$ $38$ - $40$ $44$ $64$ $66$
Average Score (Hard Tasks) 0 88 2121 - 2020 2121 $45$ $50$
@@ Appendix B Hyperparameters and Additional Details
[…]
@@
   $\mathbf{z}(t+1)\leftarrow\mathbf{z}(t)+\nicefrac{{1}}{{K}}\cdot v_{\theta}\left(\nicefrac{{t}}{{K}},\mathbf{z}(t)\leavevmode\nobreak\ |\leavevmode\nobreak\ s,a\right)$ $\triangleright$ Euler method, time $t$ is normalized
  end for
@@
   $\mathbf{x}(t+1)\leftarrow\mathbf{x}(t)+\nicefrac{{1}}{{M}}\cdot w_{\phi}\left(\nicefrac{{t}}{{M}},\mathbf{x}(t)\leavevmode\nobreak\ |\leavevmode\nobreak\ s\right)$ $\triangleright$ Euler method, time t is normalized
  end for
@@
[…]
@@
  $\mathbf{z}(0)\sim\mathrm{Unif}\leavevmode\nobreak\ [l,u],\mathbf{z}^{\prime}(0)_{1:m}\sim\mathrm{Unif}\leavevmode\nobreak\ [l,u]$ $\triangleright$ Sample initial noise for computing the Q-value
@@
[…]
@@
[…]
@@
[…]
end whilereturn One-step policy $\pi_{\omega}$
@@
Appendix C Flow Visualizations

### v2

@@ Abstract
Code and runs for floq can be found at: https://github.com/CMU-AIRe/floq
@@ 1 Introduction
Motivated by these empirical results, in this paper, we ask: can iterative computation also improve value estimation in reinforcement learning (RL)? Specifically, we are interested in improving the estimation of the Q-value function. While Q-functions map state-action inputs to a scalar value, they are known to be highly complex and difficult to fit accurately (e.g., [11]). Standard temporal-difference (TD) learning used to train Q-functions struggles to leverage capacity of deep networks [39, 40, 5, 51, 22], often resulting in poor generalization. These problems are further exacerbated in the offline RL problem setting [46, 37], where we must learn entirely from static datasets. This motivates exploring architectures that spend compute iteratively to estimate value functions, potentially yielding more accurate Q-values, and better downstream policies.
@@ 3 Background and Preliminaries
[…]
@@
[…]
@@
[…]
@@
In contrast to standard Q-functions parameterized by deep neural networks that map state-action pairs to scalar values, floq parameterizes a time-dependent, state-action-conditioned velocity field $v_{\theta}(t,\mathbf{z}\mid s,a)$ over a one-dimensional latent input $\mathbf{z}\in\mathbb{R}$. At $t=0$, this input $\mathbf{z}$ is sampled from the uniform distribution $\mathrm{Unif}\penalty 10000\ [l,u]$, where $l$ and $u$ are scalars that define the range of initial sample noise used for training. The velocity field transforms the initial sample $\mathbf{z}$ into a distribution over the Q-value. We will train floq such that the learned distribution of Q-values match a Dirac-Delta around the groundtruth Q-function, i.e., $\psi_{\theta}(1,\mathbf{z}|s,a)\sim\delta_{Q^{\pi}(s,a)}$ at $t=1$. We can obtain the Q-value sample by numerically integrating the ODE using the Euler method. One instantiation is shown below. $\forall j\leq K$:
[…]
@@ 4.2 Training Loss for the floq Architecture
[…]
@@ 4.3 Preventing Flow Collapse: How to Make floq Work Well?
Design choice 1: Distribution of the initial noise sample. As shown in prior works [48, 49], rescaling the source noise leaves the target distribution unchanged, but it alters the curvature of the transport trajectories. Interestingly, this effect seems to be particularly pronounced when applying flow-matching to scalar TD-learning (see Figure 12 in experiments). As such, we find that that setting the bounds $l$ and $u$ for the distribution of the initial noise, $\mathrm{Unif}\penalty 10000\ [l,u]$ greatly affects the performance of floq.
@@
Summary: floq architecture and training floq parameterizes the Q-function via a learned velocity field. We train this velocity field with a linear flow-matching loss against the target Q-value (Eq. 4.2). To ensure that floq scales capacity, we carefully choose the distribution of the initial noise $\mathrm{Unif}\penalty 10000\ [l,u]$ to be wide and overlapping with range of target Q-values. We encode the interpolant using a categorical representation. We represent the input $t$ to $v_{\theta}$ using a Fourier encoding. See Algorithm 1 for pseudocode and details.
@@ 5.1 Main Offline RL Results
Comparisons and evaluation protocol. In addition to a floq critic, we build our experiments to utilize flow-matching policies. This implies that flow-Q learning (FQL) [60], which utilizes a flow-matching policy but with a monolithic Q-network, is our main comparison. The FQL work reports results with 11M training steps; for a fair comparison, we additionally re-run FQL for 22M steps and report both sets of results. We run floq with a default set of hyperparameters across tasks, but on the more challenging humanoidmaze-large and antmaze-giant tasks we found a larger batch size of $512$ to be more effective, which we use for both FQL and floq on these environments. Beyond FQL, we compare against three recent state-of-the-art offline RL algorithms, all of which rely on monolithic Q-functions: (i) ReBRAC [72], which corresponds to the strongest-performing method with a monolithic Q-network and a Gaussian policy, and thus provides a competitive non-flow baseline; (ii) DSRL [74] that adapts a diffusion-based behavior cloning policy by performing RL over its latent noise space, improving over FQL; ; and (iii) SORL [12] that leverages shortcut flow models to improve the training and inference scalability of flow policies. Note that none of them utilize a flow-matching Q-function, but do innovate across various properties of policy training. Comparing to the strongest methods that innovate on the design of the policy allows us to evaluate the importance of and benefits from utilizing flows for a critic.
@@
Since DSRL [74] only evaluates on the 10 default tasks (one default task per environment), we also provide an additional results table on only the default tasks in each environment in the appendix (Table 3). On the default tasks, we find that floq outperforms DSRL (20% for DSRL vs. 45% for floq), improving by over $2\times$ in success rate. While floq uses an expected Q-value backup it still learns a stochastic Q-function. Thus, we also compare floq to a representative distributional RL approach (IQN [8]) in Table 3, and we observe that floq outperforms IQN. These results establish the efficacy of floq, and generally, performance improvements in offline RL from scaling Q-functions via flow-matching models.
@@ Acknowledgements
We thank Max Sobol Mark, Zheyuan Hu, Aravind Venugopal, Seohong Park, Mitsuhiko Nakamoto, and Amrith Setlur for feedback on an earlier version of the paper and informative discussions. We thank members of the CMU AIRe lab for feedback and support. AK thanks Max Simchowitz for discussions on flow-matching and diffusion models. This work was supported by a Schmidt Sciences AI2050 fellowship and the Office of Naval Research under N00014-24-12206. We thank the TRC program of Google Cloud and NCSA Delta for providing computational resources that supported this work. This work used the Delta advanced computing and data resource at the National Center for Supercomputing Applications (NCSA) through allocation CIS250548 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services And Support (ACCESS) program. Delta is supported by the National Science Foundation (award OAC-2005572) and the State of Illinois, and ACCESS is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
@@ Appendix A Additional Results for floq
antsoccer-arena-task2 22 $\pm 2$ 0 $\pm 1$ 96 $\pm 2$ $88$ $\pm 3$ $89$ $\pm 4$ $97$ $\pm 2$ $97$ $\pm 2$
@@
Table 3: Offline RL results (Default Tasks). floq achieves competitive or superior performance compared to the baselines. “Hard” tasks refers to the set of default tasks where the FQL baseline score is below 50% performance. floq is especially more performant on these hard tasks, more than doubling FQL’s baseline performance.
@@
Env (Default Task) BC ReBRAC DSRL SORL IQN FQL (1M) FQL(2M) floq (Def.) floq (Best)
antmaze-large 0 $\pm 0$ $91$ $\pm 10$ $40$ $\pm 29$ $93$ $\pm 2$ $86$ $\pm 1$ $80$ $\pm 8$ $85$ $\pm 4$ $94$ $\pm 4$ $94$ $\pm 4$
antmaze-giant 0 $\pm 0$ $27$ $\pm 22$ 0 $\pm 0$ $12$ $\pm 6$ 55 $\pm 6$ $11$ $\pm 16$ $14$ $\pm 29$ $70$ $\pm 8$ $86$ $\pm 4$
hmmaze-medium 11 $\pm 0$ $16$ $\pm 9$ $34$ $\pm 20$ $67$ $\pm 4$ $48$ $\pm 17$ 1919 $\pm 12$ $58$ $\pm 25$ $98$ $\pm 1$ $98$ $\pm 1$
hmmaze-large 0 $\pm 0$ 22 $\pm 1$ $10$ $\pm 12$ 2020 $\pm 9$ $35$ $\pm 5$ 88 $\pm 5$ $14$ $\pm 10$ $52$ $\pm 8$ $52$ $\pm 8$
antsoccer-arena 11 $\pm 0$ 0 $\pm 0$ $28$ $\pm 0$ $54$ $\pm 5$ $54$ $\pm 7$ $39$ $\pm 6$ $49$ $\pm 11$ $49$ $\pm 10$ $49$ $\pm 10$
cube-single 33 $\pm 1$ $92$ $\pm 4$ $93$ $\pm 14$ $99$ $\pm 0$ $98$ $\pm 1$ $96$ $\pm 1$ $94$ $\pm 5$ $99$ $\pm 2$ $99$ $\pm 2$
cube-double 0 $\pm 0$ 77 $\pm 3$ $53$ $\pm 14$ $33$ $\pm 8$ $57$ $\pm 2$ $36$ $\pm 6$ $29$ $\pm 8$ $72$ $\pm 15$ $72$ $\pm 15$
scene 11 $\pm 1$ $50$ $\pm 13$ $88$ $\pm 9$ $89$ $\pm 9$ $80$ $\pm 4$ $76$ $\pm 9$ $78$ $\pm 7$ $83$ $\pm 10$ $83$ $\pm 10$
puzzle-3x3 11 $\pm 1$ 22 $\pm 1$ 0 $\pm 0$ 2020 $\pm 3$ $16$ $\pm 5$ $14$ $\pm 4$ $17$ $\pm 6$ $17$ $\pm 6$
puzzle-4x4 0 $\pm 0$ $10$ $\pm 3$ $37$ $\pm 13$ $16$ $\pm 1$ $11$ $\pm 3$ 55 $\pm 2$ $12$ $\pm 4$ 1919 $\pm 5$
Average Score (All Tasks) 11 $30$ $38$ $49$ $40$ $44$ $64$ $66$
Average Score (Hard Tasks) 0 88 2121 $30$ 2020 2121 $45$ $50$
@@ Appendix B Hyperparameters and Additional Details
Models: a flow critic, $Q^{\text{FLOW}}_{\theta}(s,a,\mathbf{z})$, a distilled critic, $Q^{\text{distilled}}_{\psi}(s,a,\mathbf{z})$, a flow policy $\pi_{\phi}(\cdot|s)$, one-step policy $\mu_{\omega}(s,\cdot)$.
@@
    $\mathbf{z}(t+1)\leftarrow\mathbf{z}(t)+\nicefrac{{1}}{{K}}\cdot v_{\theta}\left(\nicefrac{{t}}{{K}},\mathbf{z}(t)\penalty 10000\ |\penalty 10000\ s,a\right)$ $\triangleright$ Euler method, time $t$ is normalized
@@
    $\mathbf{x}(t+1)\leftarrow\mathbf{x}(t)+\nicefrac{{1}}{{M}}\cdot w_{\phi}\left(\nicefrac{{t}}{{M}},\mathbf{x}(t)\penalty 10000\ |\penalty 10000\ s\right)$ $\triangleright$ Euler method, time t is normalized
@@
   Sample batch $\{(s,a,r,s^{\prime})\}\sim{\mathcal{D}}$
   $\triangleright$ Train vector field $v_{\theta}$ in flow critic $Q^{\text{FLOW}}_{\theta}$
@@
   $\mathbf{z}(0)\sim\mathrm{Unif}\penalty 10000\ [l,u],\mathbf{z}^{\prime}(0)_{1:m}\sim\mathrm{Unif}\penalty 10000\ [l,u]$ $\triangleright$ Sample initial noise for computing the Q-value
@@
   Update $\theta$ to minimize $\mathbb{E}\left[(v_{\color[rgb]{0.1,0.4,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.4,0.9}\theta}\left(t,\mathbf{z}(t)\penalty 10000\ |\penalty 10000\ s,a\right)-\big(\mathbf{z}(1)-\mathbf{z}(0)\big))^{2}\right]$ $\triangleright$ Linear flow-matching loss
   $\triangleright$ Train distill critic $Q^{\text{distill}}_{\psi}$ for policy extraction
   Update $\psi$ to minimize $\mathbb{E}_{\mathbf{z}(0)}\left[(Q^{\text{distill}}_{\psi}(s,a)-Q^{\text{FLOW}}_{\theta}(s,a,\mathbf{z}(0)))^{2}\right]$
   $\triangleright$ Train a BC flow policy $\pi_{\phi}$, analogous to FQL
@@
   Update $\phi$ to minimize $\mathbb{E}\left[\|w_{\color[rgb]{0.1,0.4,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.4,0.9}\phi}\left(t,\mathbf{x}(t)|s)-(\mathbf{x}(1)-\mathbf{x}(0)\right)\||_{2}^{2}\right]$ $\triangleright$ Flow-matching loss for policy
   $\triangleright$ Train one-step policy $\mu_{\omega}$ to maximize the learned distill critic while staying close to BC flow policy
@@
   Update $\omega$ to minimize $\mathbb{E}\left[-Q^{\text{distill}}_{\psi}(s,a^{\pi})+\alpha\|a^{\pi}-\pi_{\phi}(s,z)\|_{2}^{2}\right]$ return One-step policy $\pi_{\omega}$
@@
Appendix C Wall Clock Run-Time
We report the wall clock run-times for FQL and floq in Table 10.
Table 10: Total wall-clock runtime (in $10^{3}$ seconds) for FQL and floq with varying numbers of flow integration steps across four representative environments. Reported numbers correspond to 2M training steps.
Environment FQL floq (Flow Steps)
1 2 4 8 16
HM-Maze Large 14 24 28 35 50 79
HM-Maze Medium 12 19 21 23 30 47
Cube-Double 10 15 16 17 19 26
Antmaze-Giant 10 20 24 30 45 74
Appendix D Flow Visualizations